# Exploring filament galaxies using AstroSat/ UVIT

### v1

@@ 6 Impact on the mass assembly
The results presented in the previous sections establishes that the filaments provide an in-homogeneous environment for galaxies to evolve. We detect a negative gradient in FUV sSFRs on approaching the filament spine. Moreover, we find hints of gas accretion on the filament galaxies by comparing their UV and optical morphology. The observations suggest existence of physical mechanisms, such as galaxy interaction, gas starvation, inside filaments. The study of the stellar mass assembly of filament galaxies would highlight the mechanisms dominantly influencing their evolution. Multiple physical in-situ and externally driven processes may lead to stellar mass accumulation in galaxies over the cosmic time. The in-situ process includes star formation, gas recycling, and gas accretion from the inter-galactic medium (Fall & Efstathiou, 1980; Carroll & Ostlie, 2017) whereas merger-interaction, accretion of satellite galaxies, gas stripping (Toomre et al., 1977) are a few externally driven processes that steer stellar mass growth. There are two prominent mass assembly mechanisms are thought to be in action in galaxies, i.e., ‘inside-out’ and ‘outside-in’ modes in the literature (Sánchez-Blázquez et al., 2007; Pérez et al., 2013; Pan et al., 2015). In the case of inside-out mass assembly, the central regions of a galaxy form stars earlier and more rapidly than the outer regions. Outside-in growth, on the other hand, describes a pattern where star formation starts in the outer regions of a galaxy and then progresses inward toward the central region. The secularly evolving galaxies mainly grow via an inside-out mechanism whereas galaxies influenced due to environment show outside-in stellar mass build-up.
Broadband colors of galaxies are often used to determine stellar population age of galaxies in case of unavailability of spectral information. In this section, we use UV-optical () color index to indicate stellar population age. The () color is used actively to segregate star-forming and passive galaxies (Hammer et al., 2012). Both SDSS and UVIT have similar resolution and pixel scale ( 0.4). We perform photometry on a central aperture of r = 1.5 as it is roughly comparable to the point-source function full-width half maxima (PSF FWHM) of the two surveys. The radius is equivalent to 2.25 kpc at between 0.07 and 0.08. We consider the Petrosian magnitude of the sources as a measure of their total magnitude. The color gradient, , is computed using the following equation:
@@
We calculate the ratio of FUV SFR within a circular aperture of radius 1.5 to total FUV SFR. The left panel of Figure 7 shows the central-to-total SFR ratio vs. distribution for filament galaxies. Here, galaxies belonging to the blue cloud in the upper right panel of Figure 3 are termed star-forming, whereas the remaining sample detected in UVIT is grouped as quenched. The quenched galaxies in the sample have the most negative color gradients, similar to blue-cored galaxies. The star formation in these galaxies is restricted to the inner region. On the other hand, the red-cored filament galaxies have comparable FUV SFR throughout their extent.
The variation of and D is shown in the right panel of Figure 7. The sample has a nearly equal number of red- and blue-cored filament galaxies. Galaxies with M 10 M mostly have flat color gradients with few exceptions showing a blue-core. Nearly all galaxies closer to the filament spine, i.e., D 0.7 Mpc, are blue-cored regardless of their stellar masses. The result implies that these galaxies are likelier to witness galaxy nurture activities. The galaxies situated radially outwards in the filament are red-cored or, in other words, follow ‘inside-out’ mass assembly. We compare the strength of the monotonic relation between and D against and M by computing their Spearman correlation coefficient () for both pairs. The quantity D proxy for the environment ( = 0.377 with a p-value of 0.123) has a stronger positive correlation compared to the stellar mass ( = 0.226 with a p-value of 0.367). We conclude that the environment has a more substantial influence on the mass assembly of filament galaxies. However, it is essential to note that both correlations are positive and relatively weak.
@@ 7 Discussion and Conclusions
This study aims at understanding the star formation properties of galaxies inside filaments. We conduct a deep FUV survey to identify galaxies undergoing recent star formation and to subsequently investigate their FUV sSFRs, FUV morphology, and mass-assembly mechanism. The results are compared with cluster galaxies in a fixed stellar mass range. Although the filament comprises both early- and late-type galaxies, we mostly detect star-forming and disky galaxies in our FUV survey. Our UV survey does not detect galaxies closer than 0.4 Mpc from the spine. Previous studies have pointed out that the fraction of elliptical galaxies increases as the distance towards the filament axis decreases (Kuutma et al., 2017). The massive or red galaxies are more likely to exist near the spine of the filament (Chen et al., 2017). The non-detection of such galaxies in our deep UV survey confirms their dearth of star formation deep inside filaments. The mixed galaxy population type detected in filament could result from the pre-processing of galaxies inside filaments (Castignani et al., 2022).
@@
Moreover, we find a trend between the recent star formation timescale in galaxies with D. The galaxies at the filament outskirts are more star-forming than those in the inner region. The result corroborates the finding of Kuutma et al. (2017), where the spiral galaxy fraction increases with increasing distance from the spine of the filament. On comparing the observed trend with the cluster environment, the sSFR of filament galaxies increases as a function of distance to the filament axis. In contrast, cluster galaxies show no such effects. A previous study on cluster galaxies shows very slight variation wherein the sSFR of galaxies increases slightly when approaching cluster center (Laganá & Ulmer, 2017). According to a comparative study done by Mahajan et al. (2018) on filament, void, and cluster galaxies observes a similar result by analyzing their H equivalent width and colors. Ricciardelli et al. (2014) found a decrease in the SFR as distance from the void center increases up to a certain mean distance while no effect could be established for sSFRs of galaxies. The discussion above illustrates the difference in the evolution of galaxies in various LSSs. A comprehensive study combining more UV observations of LSSs would be aimed.
@@
In this study, we observe that galaxies with M than 10 M placed in the outer region of filaments have equivalent or slightly extended FUV disk size compared to their optical counterpart. The result may be an outcome of cold-mode gas accretion on the galaxies through filaments. However, some of these galaxies host AGN or are surrounded by close neighboring galaxies, which may have led to the observed result. Therefore, our study could neither confirm nor deny the cold mode accretion process reported by Kleiner et al. (2017). Agreeably, we detected a limited number of filament galaxies detected in our UVIT survey. Hence, we aim to perform a similar study with an increased sample to enhance our understanding of filament galaxies. Deep multi-band observations could be utilized to undermine the prevalence cold-mode gas accretion in filaments.
@@
We study the mass assembly of filament galaxies based on color gradient. The galaxies closer to the filament spine mainly evolve via outside-in growth mass assembly. In other words, galaxies host blue cores closer to the filament axis. Similar results were reported by Lee et al. (2021), where the distribution of H equivalent width as a function of vertical distance from the filament axis was studied. The author showed that the high stellar mass galaxies closer to the filament spine are actively star-forming in their central region. A considerable cause of the central star formation could be the increased likelihood of merger interactions.
@@
Our work establishes that filament could serve as a hotbed for several activities impacting the growth of galaxies residing inside filaments., e.g., galaxy interactions, gas starvation, and gas accretion. These processes attribute to a range of galaxy properties in color, SFR, and morphology.
The results from our study are summarized below:
i) We detect galaxies with D 0.4 Mpc in our deep UV survey.
@@
ii) The study establishes that the filament environment substantially affects the evolution of galaxies inside them. We find early- and late-type galaxies residing inside the filament. Most of the galaxies detected by in our FUV survey are star-forming and late-type.
@@
iii) There is a positive gradient in FUV sSFR as the distance from the filament axis increases.
iv) We find that massive galaxies in the filament are undergoing recent star formation on a rapid timescale compared to galaxies in clusters.
@@
v) The UV morphology for a few galaxies residing in the outer edge of filaments is extended more than their optical profile.
@@
vi) ‘Outside-in’ mass-assembly mode is more prevalent in the inner region of the filaments, i.e., galaxies inside filaments are centrally star-forming. The result could be due to an increased rate of galaxy merger interaction in the inner region.

### v2

@@ 6 Impact on the mass assembly
The results presented in the earlier sections establish that the filaments provide an in-homogeneous environment for galaxies to evolve. We detect negative gradients in FUV sSFRs on approaching the filament spine. Moreover, we find hints of gas accretion on the filament galaxies by comparing their UV and optical morphology. The observations suggest the existence of physical mechanisms, such as galaxy interaction and gas starvation deep inside filaments. The study of the stellar mass assembly of filament galaxies would highlight the mechanisms dominantly influencing their evolution. Multiple physical in-situ and externally driven processes may lead to stellar mass accumulation in galaxies over the cosmic time. The in-situ process includes star formation, gas recycling, and gas accretion from the inter-galactic medium (Fall & Efstathiou, 1980; Carroll & Ostlie, 2017) whereas merger-interaction, accretion of satellite galaxies, gas stripping (Toomre et al., 1977) are a few externally driven processes that steer stellar mass growth. Two prominent mass assembly mechanisms are considered in action in galaxies, i.e., ‘inside-out’ and ‘outside-in’ modes in the literature (Sánchez-Blázquez et al., 2007; Pérez et al., 2013; Pan et al., 2015). In the case of inside-out mass assembly, the central regions of a galaxy form stars earlier and more rapidly than the outer regions. Outside-in growth, on the other hand, describes a pattern where star formation starts in the outer regions of a galaxy and then progresses inward toward the central region. The secularly evolving galaxies mainly grow via an inside-out mechanism, whereas galaxies influenced by the external environment show outside-in stellar mass build-up.
There are many excellent probes to investigate the stellar mass assembly of galaxies, e.g., a combination of well-resolved near-IR and H observations which trace long-term and very recent ( 10 Myr) star formation, respectively, would be effective in tracing the influence of gas accretion and longer-term inside-out star formation history. However, we lack such observations for our sample. Broadband colors of galaxies are another alternative often used to determine the stellar population age of galaxies. This section uses the UV-optical () color index to indicate stellar population age. The () color is used actively to segregate star-forming and passive galaxies (Hammer et al., 2012). Both SDSS and UVIT have similar resolution and pixel scale ( 0.′′4). We perform photometry on a central aperture of r = 1.5′′ as it is roughly comparable to the point-source function full-width half maxima (PSF FWHM) of the two surveys. The radius is equivalent to 2.25 kpc at between 0.07 and 0.08. Note that AGNs are excluded in this analysis.
We consider the Petrosian magnitude of the sources as a measure of their total magnitude. The color gradient, , is computed using the following equation:
@@
We calculate the ratio of FUV SFR within a circular aperture of radius 1.′′5 to total FUV SFR. The left panel of Figure 7 shows the central-to-total SFR ratio vs. distribution for filament galaxies. Here, galaxies belonging to the blue cloud in the upper right panel of Figure 3 are termed star-forming, whereas the remaining sample detected in UVIT is grouped as quenched. The quenched galaxy in the sample has the most negative color gradient, similar to blue-cored galaxies. The star formation in such galaxies is restricted to the inner region. On the other hand, the red-cored filament galaxies have comparable FUV SFR throughout their extent.
The variation of and Dfil is shown in the right panel of Figure 7. Galaxies with M⋆ 1010.5 M⊙ mostly have flat color gradients with few exceptions showing a blue-core. A majority of blue-cored galaxies lie closer to the filament spine. The result implies that these galaxies are likelier to witness galaxy nurture activities. The galaxies situated radially outwards in the filament are red-cored or, in other words, follow ‘inside-out’ mass assembly. We compare the strength of the monotonic relation between and Dfil against and M⋆ by computing their Spearman correlation coefficient () for both pairs. We find that the quantities do not correlate significantly with .
@@ 7 Discussion and Conclusions
This study aims at understanding the star formation properties of galaxies inside filaments. We conduct a deep FUV survey to identify galaxies undergoing recent star formation and to subsequently investigate their FUV sSFRs, FUV morphology, and mass-assembly mechanism. The results are compared with cluster galaxies in a fixed stellar mass range. Although the filament comprises both early- and late-type galaxies, we mostly detect star-forming and disky galaxies in our FUV survey. Our UV survey does not detect galaxies closer than 0.4 Mpc from the spine. Previous studies have pointed out that massive or red galaxies are more likely to exist near the spine of the filament (Chen et al., 2017; Kuutma et al., 2017). The non-detection of such galaxies in our deep UV survey confirms the lack of star formation deep inside filaments. The mixed galaxy population type detected in filament could result from the pre-processing of galaxies inside filaments (Castignani et al., 2022).
@@
Moreover, we find a positive trend between sSFRFUV in galaxies with Dfil. The galaxies at the filament outskirts are more star-forming than those in the inner region. The result corroborates the finding of Kuutma et al. (2017), where the spiral galaxy fraction increases with increasing distance from the spine of the filament. The cluster galaxies show no variation SFR (or sSFR) across the cluster environment (Laganá & Ulmer, 2017). However, some studies report a drop in SFR with cluster-centric distances (Balogh et al., 2004; Mahajan et al., 2018). In the case of void galaxies, Ricciardelli et al. (2014) found a decrease in the SFR as the distance from the void center increases to a certain mean distance, while no effect could be established for the sSFRs of galaxies. The discussion above illustrates the difference in the evolution of galaxies in various LSSs. A comprehensive study combining more UV observations of LSSs would be aimed in the future.
@@
In this study, we observe that a few intermediate-to-high stellar mass galaxies residing in the outer region of filaments have equivalent or slightly extended FUV disk sizes compared to their optical counterpart. The result may be an outcome of cold-mode gas accretion on the galaxies through filaments. However, some of these galaxies host AGN or are surrounded by close neighboring galaxies, which may have led to the observed result. Therefore, our study could neither confirm nor deny the cold mode accretion process reported by Kleiner et al. (2017). Agreeably, we detected a limited number of filament galaxies in our UVIT survey. Hence, we aim to perform a similar study with an increased sample to enhance our understanding of filament galaxies. Deep multi-band observations could be utilized to uncover the prevalence of cold-mode gas accretion in filaments.
@@
We study the mass assembly of filament galaxies based on color gradient. The galaxies closer to the filament spine mainly evolve via outside-in growth mass assembly. Similar results were reported by Lee et al. (2021), where the high stellar mass galaxies closer to the filament spine show elevated Hα equivalent widths, indicating central star formation. A considerable cause of the central star formation could be the increased likelihood of merger interactions. Our results largely depend on the estimated Dfil, which may represent the lower limit of the actual values due to the projection effect.
@@
Our work establishes that filament could serve as a hotbed for several activities impacting the growth of galaxies residing inside filaments., e.g., galaxy interactions, gas starvation, and gas accretion. As a result, the filament galaxies display a wide range of galaxy properties in color, SFR, and morphology. The results from our study are summarized below:
i) We do not detect galaxies within Dfil 0.4 Mpc in our deep UV survey.
@@
ii) The study establishes that the filament environment substantially affects the evolution of galaxies inside them. We find early- and late-type galaxies residing inside the filament. Most of the galaxies detected in our FUV survey are star-forming and late-type.
iii) The FUV SFR for star-forming filament galaxies is in agreement with SFR144MHz.
@@
iv) We witness an increase in FUV sSFR as the distance from the filament axis increases.
v) The filament galaxies show higher sSFRFUV compared to cluster galaxies in a fixed same stellar mass range.
@@
vi) The UV morphology for a few galaxies residing in the outer edge of filaments is extended more than their optical profile.
@@
vii) ‘Inside-out’ mass-assembly mode is more prevalent in the outer region of the filaments than in the inner region of the filament. The result could be due to an increased rate of galaxy merger interaction closer to the filament spine.